You're dressing in the dark because of a power failure. You have four black socks and eight red socks on a drawer. You pull out three socks. (Hypergeometric Probability Distribution)

a) What is the probability that two are black? Carry answer to the nearest ten-thousandth
Answer #1

N = 4 + 8 = 12

n = 3

k = 4

a) P(X = 2) = 4C2 * 8C1 / 12C3 = 0.2182